# Automated Detection of Galactic Rings from SDSS Images

### v1

@@ Abstract
Morphological features in galaxies, like spiral arms, bars, rings, tidal tails etc. carry information about their structure, origin and evolution. It is therefore important to catalogue and study such features and to correlate them with other basic galaxy properties the environment in which the galaxies are located and their interactions with other galaxies. Surveys such as SDSS, Pan-STARRS, HSC-SSP have made available very large samples of galaxies for gainful morphological studies. The availability of galaxy images and catalogues will increase manifold with future surveys like LSST. The volume of present and future data is so large that traditional methods, which involve expert astronomers identifying morphological features through visual inspection, are no longer sufficient. It is therefore necessary to use AI based techniques like machine learning and deep learning for finding morphological structures quickly and efficiently. We report in this study the application of deep learning for finding ring like structures in galaxy images from the Sloan Digital Sky Survey (SDSS) data release DR18. We use a catalogue by Buta (2017) of ringed galaxies from the SDSS to train the network reaching good accuracy and recall, and generate a catalogue of 29420 galaxies of which 9805 have ring like structures with prediction confidence exceeding 90 percent. Using a catalogue of barred galaxy images identified by Abraham et. al. (2018) using deep learning techniques, we identify a set of 2087 galaxies with bars as well as rings. The catalogues should be very useful in understanding the origin of these important morphological structures.
@@ 1 Introduction
In recent decades, the availability of many millions of higher resolution multiwavelength images from large deep field surveys made using modern instrumentation have made possible detailed morphological studies (Buta, 2011). Galaxy morphology plays a key role in bettering our understanding of the secular processes that underlie galaxy evolution. Several important questions concerning the formation and evolution of galaxies can be addressed from the observed change in morphology with epoch.
A fairly common feature of disk galaxies is the presence of a ring shaped pattern in their light distribution (Buta & Combes, 1996). These rings in galaxies have come to be considered an integral part of galaxy morphology (Kormendy, 1979). Recognizing their importance, the de Vaucouleurs Revised Hubble-Sandage Classification System (VRHS; De Vaucouleurs, 1959) added rings as another dimension to the two-dimensional Hubble tuning fork, turning it into a classification volume. Rings in galaxies can fundamentally be divided into normal rings, also called ‘resonance rings’ and ‘catastrophic rings’, which are a result of galaxy collisions (Buta & Combes, 1996). Although different subclasses exist within the catastrophic rings, such as polar, accretion, and collisional rings, they constitute a very small minority of all observed rings. Madore et al. (2009) have estimated the abundance of collisional rings to be only 1 in 1000. Several theories have been put forward to explain the different ring structures that are observed, with the resonance interpretation being the most popular. The “Manifold Theory” that has been proposed recently Athanassoula et al. (2009) also has had considerable success in explaining several aspects of ring morphology.
Several catalogues that contain the visual morphology classification of galaxies are available today. Notable among them are Nair & Abraham (2010) and Fukugita et al. (2007), with the former containing the detailed classifications of 14,034 galaxies. Crowd-sourcing efforts such as the Galaxy Zoo citizen science project (Lintott et al., 2011) have managed to ramp up the number of galaxy morphology classifications to nearly 900,000 by involving volunteers. This vast database has motivated others to pick specific subtypes of galaxies for detailed classifications. For example, Buta (2017) has exhaustively classified 3962 ringed galaxies taken from this set. However, all these efforts involve a human in the loop for classifying galaxy morphologies.
@@
In the present study we investigate the efficacy of an automated method based on Convolutional Neural Networks for the detection of rings in galaxies. We use galaxy images from the Sloan Digital Sky Survey (SDSS; York et al., 2000) for the training and testing of our network. We treat the problem as a binary classification one, treating all sub-classes of rings as a single ring type and distinguishing it from galaxies that do not contain any rings, the non-ring type. The trained network can be used to identify ringed galaxies from the SDSS.
The rest of the paper is organized as follows: In Section 2, we describe the data that we used. The neural network architecture is introduced in Section 3. Sections 4 and 5 describe the data augmentation techniques and the training procedure for our neural network respectively. Section 6 analyses our results on the trained network. Section 7 presents the catalog of ringed galaxies generated using our trained network. Section 8 presents a subset of the ringed galaxies that have also been identified to have bars. Finally Section 9 gives the concluding remarks.
@@ 2 Data
The performance of a supervised machine learning classifier depends to a large extent on the quality of the labelled data that is made available to it. For deep learning models the quantity becomes important as well because of the large number of trainable or free parameters in the model. Many of the standard datasets available for galaxy morphology classification do not contains rings as a class (see, e.g., DECALS10  [^1] https://astro.utoronto.ca/~hleung/shared/Galaxy10/Galaxy10_DECals.h5). Even with the Galaxy Zoo 2 dataset (Willett et al., 2013), ringed galaxies for which at least 50 percent of zoo volunteers agreed on a response are relatively small in number.
@@ 2.1 Rings and Non-Rings
We used the Buta (2017) catalog as our primary source to identify galaxies with rings. The author has provided detailed visual morphological classification of 3962 galaxies with rings. The classifications are done within the framework of the Comprehensive de Vaucouleurs revised Hubble-Sandage system (CVRHS; Buta et al., 2015, 2007). The catalog also provides the author’s comments that are useful for creating a good quality training set by identifying rare cases of rings, poorly resolved galaxies etc. The galaxies used for classification by the author were picked from the Galaxy Zoo 2 (Willett et al., 2013) survey (hereafter, GZ2).
Along with a good quality training sample for rings, it is equally important to obtain a good sample of galaxies that do not have any presence of rings. Since the Buta (2017) catalog does not contain galaxies without rings, we used the Nair & Abraham (2010) catalog, which is one of the largest catalogs for the visual morphological classification of galaxies. They used galaxy images from the SDSS DR4 release (Stoughton et al. (2002); Adelman-McCarthy et al. (2006)). Galaxies with spectroscopic redshifts in the range $0.10.01<z<$ 0.1 and extinction corrected g’-band magnitude $<16$ were selected from the spectroscopic main sample (Strauss et al., 2002). This led to their final sample size of 14,038 galaxies. Figure 1 shows the actual distribution of redshift and extinction corrected g-magnitude for the ringed and non-ringed galaxies in our training set.
Figure 1: The normalized distribution of redshifts and extinction corrected g-mag for the ringed galaxies(left) and non-ringed galaxies(right)
The classification scheme used in Nair & Abraham (2010) is not the CVRHS. Rather it is primarily based on the Carnegie Atlas of Galaxies (Sandage & Bedke, 1994) used in consultation with the Third Reference Catalog of Bright Galaxies (RC3; de Vaucouleurs et al., 1991). In order to use only galaxies without rings, we selected from the catalog galaxies that have the ring type and ring flag column set to zero. Along with ringed galaxies identified by Buta (2017), we also used the ringed galaxies identified by Nair & Abraham (2010). Their catalog has identified 2598 rings from a total of 14034 galaxies. By visual inspection we identified 800 galaxies that had clearly identifiable rings and added this to our training sample.
@@ 2.2 Selection Criteria
The images used for the GZ2 survey come from the SDSS Data Release 7 (DR7; Abazajian et al., 2009). The GZ2 includes galaxies with extinction corrected Petrosian half-light magnitude in the r-band < $17.0$, Petrosian radius $\tt{petroR90_{r}}>3$ arcsec and spectroscopic redshift in the range $0.250.0005<z<$ 0.25, when it is known. Finally, galaxies that have SDSS flags which are either SATURATED, BRIGHT, OR BLENDED without an accompanying NODEBLEND were removed. We further selected galaxies that have an extinction corrected g’-band magnitude < 16 based on the selection criteria used for the non-ring galaxies. The spectroscopic redshift was also limited to be in the same range as that of the non-ring galaxies, specifically $0.10.01<z<$ 0.1.
@@ 2.3 SDSS Image Cutouts
Most imaging surveys provide an Application Programming Interface (API) access for downloading image cutouts in bulk from their servers. For our purposes we needed to download images in bulk both for creating the training data as well for catalog preparation. We modified a python library [^2] https://pypi.org/project/panstamps/ that already exists in the public domain for downloading images. We incorporated multi-threaded parallel downloading using the Concurrent  [^3] https://docs.python.org/3/library/concurrent.futures.html python module into the script. This enabled us to download images in parallel, speeding up the whole data downloading process. The script takes as input a comma-separated-value (CSV) file containing a minimum of three columns, the Right Ascension (RA), Declination (DEC) and a unique key that can be used to name the downloaded images. Figure 2 shows a sample of the galaxies with and without rings in our training set.
Refer to caption Galaxies with rings
Refer to caption Galaxies with no rings
@@ 3 Network Design
AlexNet (Krizhevsky et al., 2012) is one of the earliest CNN architectures desgined and yet one of the most reliable. Having a smaller size than most modern architectures results in lesser training time and also the requirement of lesser amount of training data. The original AlexNet architecture was designed for the ImageNet challenge (Deng et al., 2009). The network has eight layers with the first five being convolutional layers and the remaining three being fully-connected. The first convolutional layer filters the 224 × 224 × 3 input image with 96 kernels of size 11 × 11 × 3 with a stride of four pixels. All of the convolution and fully connected layers are immediately followed by a block consisting of a Rectified Linear Unit activation (ReLU; Nair & Hinton 2010) and a batch normalisation (NORM) layer. The first, second and final convolution layers are each followed by a maxpooling layer, that leads to 50 percent downsampling of the convolution feature maps. These layers also contribute to the rotational invariance of the learnt model weights (Boureau et al., 2010). The max-pooling layers of the first two convolutions are followed by dropout layers (Srivastava et al., 2014) having a dropout fraction of 0.25. These prevent overfitting of the model by randomly dropping 25 percent of the weights during each training iteration.
We left the input layer size unchanged since it was suitable for our purposes. However, galactic rings are features that are harder to distinguish in comparison to the classes in the ImageNet challenge. Hence we used a smaller size of five for the kernels in the first convolutional layer. This seemed to have a considerable impact on the accuracy of the network. Thus in our network the first convolution layer has 96 kernels of size $5\times 5\times 3$ and the output is fed to the second layer of convolution with 256 kernels each of size $5\times 5\times 48$ The third, fourth and fifth convolutional layers have respectively 384, 384 and 256 kernels of size $3\times 3\times 256$, $3\times 3\times 192$ and $3\times 3\times 192$ respectively. These learn more complex features whose final output is fed into another max-pooling layer. The output of the final pooling layer is then fed into a series of fully connected layers of size 4096. These layers are connected by dropout layers which have 0.50 dropout fraction. The final layer is the sigmoid activation function layer which outputs a value that can be interpreted as the probability of the input belonging to the positive class. The exact architecture of our network is shown as a table and schematic diagram in Figure 3.
@@ 3.1 Selection of Hyperparameters
Hyperparameter tuning is an important part of any machine learning or deep learning project. Hyperparameters are those parameters that are not learnt or updated during the learning process and hence need to be manually set by the programmer. These can include the network architecture itself and several other parameters like the batch size used for the stochastic gradient descent or the number of epochs of training etc. These parameters have a direct bearing on the performance and accuracy of the model. Finding the right set of hyperparameters is usually a trial and error process.
Since a machine learning model learns iteratively by making errors the loss function is an important hyperparameter to tune. The loss function specifies what constitutes an error and how much weightage should be given to individual errors. We have used categorical crossentropy [^4] https://www.tensorflow.org/api_docs/python/tf/keras/metrics/categorical_crossentropy, which is a standard loss function suitable for binary classification problems. The optimizer function has an effect on how the network weights are updated in order to minimize this loss. We used the Adam optimizer with a learning rate of 0.003. Regularization techniques are a set of safeguard measures that prevent the model from overfitting on the training data and thus losing its ability to generalize. We have used L2 weight regularization with a factor of 0.0002 to penalize the weights from growing too much during training. The batch size controls how many input samples are considered together for evaluating the loss function. This leads to a trade-off between accuracy and speed. A low batch size leads to better accuracy often at the cost of very large training times. A batch size of 16 was seen to be sufficient in our case. The number of epochs or (num epochs) decides how many times the network gets to see the complete training data. Early Stopping (Prechelt, 1998) is a technique that can be used to stop the training process if subsquent trainings do not improve the metrics we are interested in like the validation loss. However, using it can lead to uncertaininity as it introduces another hyperparameter called the patience that needs to be properly set. We trained the network for num epochs without early stopping.
@@ 4 Image augmentation
Neural networks are prone to overfitting if the data available for training is insufficient. The size of data required increases with the complexity of the network which in turn is dependent on the number of free parameters available for tuning. Increasing the training data is limited by the availability of dedicated catalogs that can provide a sufficiently large number of ringed galaxies. A standard practice in such cases is to create more examples for training the network through data augmentation. This involves the transformation of existing images to generate realistic looking new images. CNNs are translation invariant by design but they are not rotation invariant. Additionally, individual CNN filters although translational invariant are not invariant to rotations of the images. However ringed galaxies have rotational symmetry and hence their classification should be invariant to rotations. By transforming each of the training images using random rotations we could increase the dataset size at the same time making the network invariant to rotations. We have used operations such as horizontal and vertical flipping, rotation through arbitrary angles, brightness and contrast adjustments for augmenting our training set. The augmentations that we have used are illustrated in Figure 4 using a random galaxy.
We implemented image augmentations using the on-the-fly mode provided by the keras-tensorflow (Chollet et al., 2015) framework . In this way, during training, each image from the training set is taken and a predefined sequence of transformations are applied to it. The parameters used for the transformations are randomized so that each epoch, the network sees a different set of augmented images. In this way we avoid having to generate all the augmented images before the start of the training and also save a lot of disk space. Additionally by making use of such a data pipeline we can make the process of k-fold splitting of the dataset easier to manage.
Refer to caption Original
Refer to caption Horizontal Flip
Refer to caption Vertical Flip
Refer to caption Rotation
Refer to caption Brightness
Refer to caption Contrast

### v2

@@ Abstract
Morphological features in galaxies, like spiral arms, bars, rings, tidal tails etc. carry information about their structure, origin and evolution. It is therefore important to catalog and study such features and to correlate them with other basic galaxy properties, the environment in which the galaxies are located and their interactions with other galaxies. The volume of present and future data on galaxies is so large that traditional methods, which involve expert astronomers identifying morphological features through visual inspection, are no longer sufficient. It is therefore necessary to use AI based techniques like machine learning and deep learning for finding morphological structures quickly and efficiently. We report in this study the application of deep learning for finding ring like structures in galaxy images from the Sloan Digital Sky Survey (SDSS) data release DR18. We use a catalog by Buta (2017) of ringed galaxies from the SDSS to train the network, reaching good accuracy and recall, and generate a catalog of 29420 galaxies of which 4855 have ring like structures with prediction confidence exceeding 90 percent. Using a catalog of barred galaxy images identified by Abraham et. al. (2018) using deep learning techniques, we identify a set of 2087 galaxies with bars as well as rings. The catalog should be very useful in understanding the origin of these important morphological structures. As an example of the usefulness of the catalog, we explore the environments and star formation characteristics of ring galaxies in our sample.
@@ 1 Introduction
In recent decades, the availability of many millions of higher resolution multiwavelength images from large deep field surveys with modern instrumentation have made possible detailed morphological studies (Buta, 2011). Galaxy morphology plays a key role in bettering our understanding of the secular processes that underlie galaxy evolution. Several important questions concerning the formation and evolution of galaxies can be addressed from the observed change in morphology with epoch.
A fairly common feature of disk galaxies is the presence of a ring shaped pattern in their light distribution (Buta & Combes, 1996). These rings in galaxies have come to be considered an integral part of galaxy morphology (Kormendy, 1979). Recognizing their importance, the de Vaucouleurs Revised Hubble-Sandage Classification System (VRHS; De Vaucouleurs, 1959) added rings as another dimension to the two-dimensional Hubble tuning fork, turning it into a classification volume. Rings in galaxies can fundamentally be divided into normal rings, also called ‘resonance rings’ and ‘catastrophic rings’, which are a result of galaxy collisions (Buta & Combes, 1996). Although different sub-classes exist within the catastrophic rings, such as polar, accretion, and collisional rings, they constitute a very small minority of all observed rings. Madore et al. (2009) have estimated the abundance of collisional rings to be only about 1 in 1000. Several theories have been put forward to explain the different ring structures that are observed, with the resonance interpretation being the most popular. The “Manifold Theory” that has been proposed Athanassoula et al. (2009) also has had considerable success in explaining several aspects of ring morphology.
Several catalogs that contain the visual morphology classification of galaxies are available today. Notable among them are Nair & Abraham (2010) and Fukugita et al. (2007), with the former containing the detailed classifications of 14,034 galaxies. Crowd-sourcing efforts such as the Galaxy Zoo citizen science project (Lintott et al., 2011) have managed to ramp up the number of galaxy morphology classifications to nearly 900,000 by involving volunteers. This vast database has motivated others to pick specific subtypes of galaxies for detailed classifications. For example, Buta (2017) has exhaustively classified 3962 ringed galaxies taken from this set. However, all these efforts involve a human in the loop for classifying galaxy morphologies.
@@
In the present study we investigate the efficacy of an automated method based on Convolutional Neural Networks for the detection of rings in galaxies. We use galaxy images from the Sloan Digital Sky Survey (SDSS; York et al., 2000) for the training and testing of our network. We treat the problem as a binary classification one, treating all sub-classes of rings as a single ring type and distinguishing it from galaxies that do not contain any rings, the non-ring type. The trained network can be used to identify ringed galaxies from the SDSS. We used the network’s predictions to generate a sample of 4855 ringed galaxies with a classification threshold of 0.90. Using this sample we explored the connection between ring formation and star formation activity considering the main sequence, green valley and quenched galaxy populations. We also analyzed the star formation by examining their distribution across different environments, categorized by their local surface densities.
The rest of the paper is organized as follows: In Section 2 we describe the data used in the analysis and introduce the neural network architecture in Section 3. We describe the data augmentation techniques and training procedure for the neural network in Sections 4 and 5 respectively and in Section 6 we analyse the results on the trained network. In Section 7 we present the catalog of ringed galaxies generated using our trained network and in Section 8 we present a subset of the ringed galaxies that have also been identified to have bars. In Section 9 we provide a discussion regarding the environments and star formation rates for the ringed galaxies in our catalog. Finally in Section 10 we give the concluding remarks.
@@ 2 Data
The performance of a supervised machine learning classifier depends to a large extent on the quality of the labelled data that is made available to it. For deep learning models the quantity becomes important as well because of the large number of trainable or free parameters in the model. Many of the standard datasets available for galaxy morphology classification do not contain rings as a class (see, e.g., DECALS10  [^1] https://astro.utoronto.ca/~hleung/shared/Galaxy10/Galaxy10_DECals.h5). Even with the Galaxy Zoo 2 dataset (Willett et al., 2013), ringed galaxies for which at least 50 percent of zoo volunteers agreed on a response are relatively small in number.
@@ 2.1 Rings and Non-Rings
We used the Buta (2017) catalog as our primary source to identify galaxies with rings. The author has provided detailed visual morphological classification of 3962 galaxies such galaxies. The classifications are done within the framework of the comprehensive de Vaucouleurs revised Hubble-Sandage system (CVRHS; Buta et al., 2015, 2007). The catalog also provides the author’s comments that are useful for creating a good quality training set by identifying rare cases of rings, poorly resolved galaxies etc. The galaxies used for classification by the author were picked from the Galaxy Zoo 2 (Willett et al., 2013) survey (hereafter, GZ2).
Along with a good quality training sample for rings, it is equally important to obtain a good sample of galaxies that do not have any presence of rings. Since the Buta (2017) catalog does not contain such galaxies, we used the Nair & Abraham (2010) catalog, which is one of the largest catalogs for the visual morphological classification of galaxies. They used galaxy images from the SDSS DR4 release (Stoughton et al. (2002); Adelman-McCarthy et al. (2006)). Galaxies with spectroscopic redshifts in the range $0.10.01<z<$ 0.1 and extinction corrected g’-band magnitude $<16$ were selected from the spectroscopic main sample (Strauss et al., 2002). This led to their final sample size of 14,034 galaxies.
Figure 1: The normalized distribution of redshifts and extinction corrected g-magnitude for the ringed galaxies(left) and non-ringed galaxies(right)
The classification scheme used in Nair & Abraham (2010) is not the CVRHS. Rather it is primarily based on the Carnegie Atlas of Galaxies (Sandage & Bedke, 1994) used in consultation with the Third Reference Catalog of Bright Galaxies (RC3; de Vaucouleurs et al., 1991). In order to use only galaxies without rings, we selected from the catalog galaxies that have the ring type and ring flag column set to zero.
@@ 2.2 Selection Criteria
The images used for the GZ2 survey come from the SDSS Data Release 7 (DR7; Abazajian et al., 2009). The GZ2 includes galaxies with extinction corrected Petrosian half-light magnitude in the r-band < $17.0$, Petrosian radius $\tt{petroR90_{r}}>3$ arcsec and spectroscopic redshift in the range $0.250.0005<z<$ 0.25, when it is known. Finally, galaxies that have SDSS flags which are either SATURATED, BRIGHT, OR BLENDED without an accompanying NODEBLEND were removed. We further selected galaxies that have an extinction corrected g’-band magnitude < 16 based on the selection criteria used for the non-ring galaxies. The spectroscopic redshift was also limited to be in the same range as that of the non-ring galaxies, specifically $0.10.01<z<$ 0.1. By visual inspection we removed 37 ringed and 797 non-ringed galaxies which contained artefacts and those which were not clearly distinguishable. After the selection criteria was applied and the manual removal, the number of ringed galaxies obtained were 1122 and the number of galaxies without rings came out to be 10,639. Figure 1 shows the distribution of redshift and extinction corrected g-magnitude for the ringed and non-ringed galaxies in our training set.
@@ 2.3 SDSS Image Cutouts
We modified a python library [^2] https://pypi.org/project/panstamps/ that exists in the public domain for downloading images we needed for training as well as catalog preparation. We incorporated multi-threaded parallel downloading using the Concurrent  [^3] https://docs.python.org/3/library/concurrent.futures.html python module into the script to speed up the downloading process. Figure 2 shows a sample of the galaxies with and without rings in our training set.
Galaxies with rings
Galaxies with no rings
@@ 3 Network Design
AlexNet (Krizhevsky et al., 2012) is one of the earliest CNN architectures designed and yet one of the most reliable. Having a smaller size than most modern architectures results in lesser training time and also the requirement of lesser amount of training data for this network. The original AlexNet architecture was designed for the ImageNet challenge (Deng et al., 2009). The network has eight layers with the first five being convolutional layers and the remaining three being fully-connected. The first convolutional layer filters the 224 × 224 × 3 input image with 96 kernels of size 11 × 11 × 3 with a stride of four pixels. All of the convolution and fully connected layers are immediately followed by a block consisting of a Rectified Linear Unit activation (ReLU; Nair & Hinton 2010) and a batch normalisation (NORM) layer. The first, second and final convolution layers are each followed by a maxpooling layer, that leads to 50 percent downsampling of the convolution feature maps. These layers also contribute to the rotational invariance of the learnt model weights (Boureau et al., 2010). The max-pooling layers of the first two convolutions are followed by dropout layers (Srivastava et al., 2014) having a dropout fraction of 0.25. These prevent overfitting of the model by randomly dropping 25 percent of the weights during each training iteration.
We left the input layer size unchanged since it was suitable for our purposes. However, galactic rings are features that are harder to distinguish in comparison to the classes in the ImageNet challenge. We therefore used a smaller size of five for the kernels in the first convolutional layer. This seemed to have a considerable impact on the accuracy of the network. Thus in our network the first convolution layer has 96 kernels of size $5\times 5\times 3$ and the output is fed to the second layer of convolution with 256 kernels each of size $5\times 5\times 48$ The third, fourth and fifth convolutional layers have respectively 384, 384 and 256 kernels of size $3\times 3\times 256$, $3\times 3\times 192$ and $3\times 3\times 192$ respectively. These learn more complex features and their final output is fed into another max-pooling layer. The output of the final pooling layer is then fed into a series of fully connected layers of size 4096. These layers are connected by dropout layers which have 0.50 dropout fraction. The final layer is the sigmoid activation function layer which outputs a value that can be interpreted as the probability of the input belonging to the positive class. The architecture of our network is shown as a table and schematic diagram in Figure 3.
@@ 3.1 Selection of Hyperparameters
Hyperparameters are those parameters that are not learnt or updated during the learning process and hence need to be manually set by the programmer. These can include the network architecture itself and several other parameters like the batch size used for the stochastic gradient descent or the number of epochs of training etc. Hyperparameter tuning is an important part of a deep learning project, since these parameters have a direct bearing on the performance and accuracy of the model.
The loss function, which specifies the error and how much weightage should be given to individual errors is an important hyperparameter to tune. We have used categorical crossentropy [^4] https://www.tensorflow.org/api_docs/python/tf/keras/metrics/categorical_crossentropy, which is a standard loss function suitable for binary classification problems. The optimizer function has an effect on how the network weights are updated in order to minimize this loss. We used the Adam optimizer with a learning rate of 0.003. Regularization techniques are a set of safeguard measures that prevent the model from overfitting on the training data and thus losing its ability to generalize. We have used L2 weight regularization with a factor of 0.0002 to penalize the weights from growing too much during training. The batch size controls how many input samples are considered together for evaluating the loss function. This leads to a trade-off between accuracy and speed. A batch size of 16 was seen to be sufficient in our case. The number of epochs or (num epochs) decides how many times the network gets to see the complete training data. We trained the network for num epochs without early stopping.
@@ 4 Image augmentation
Neural networks are prone to overfitting if the data available for training is insufficient. In our case the possibility of increasing the training data is limited by the availability of catalogs of ringed galaxies. It is therefore necessary to increase the training sample size, through data augmentation by transforming the available images. We have used operations such as horizontal and vertical flipping, rotation through arbitrary angles, brightness and contrast adjustments for augmenting our training set. Using rotations also makes the network invariant to rotations, which is needed because of the rotational symmetry of the ringed galaxies, The augmentations that we have used are illustrated in Figure 4 using a random galaxy.
We implemented image augmentations using the on-the-fly mode provided by the keras-tensorflow (Chollet et al., 2015) framework. Here a predefined sequence of transformations is applied to each image from the training set, using parameters which are randomized so that at each epoch, the network sees a different set of augmented images. In this way we avoid generating all the augmented images before the start of the training and also save disk space.
Original
Horizontal Flip
Vertical Flip
Rotation
Brightness
Contrast